The difference between two numbers is 7. Six times the smaller plus the larger is 77. Find the numbers.
Answer by TimothyLamb(4379)

x = larger number
y = smaller number
---
x - y = 7
x + 6y = 77
---
put the system of linear equations into standard form
---
x - y = 7
x + 6y = 77
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = larger number = 17
y = smaller number = 10
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---